You have sampled 25 students to find the mean SAT scores at Morris Knolls High School. A 95% confidence interval for the mean SAT score is 900 to 1100. Which of the following statement gives a valid interpretation of this interval?
A. 95% of the 25 students have a mean score between 900 and 1100.
B. 95% of the population of all students at Morris Knolls have a score between 900 and 1,100.
C. If this procedure were repeated many times, 95% of the resulting confidence intervals would contain the true mean SAT score at Morris Knolls.
D. If this procedure were repeated many times, 95% of the sample means would be between 900 and 1100.
E. If 100 samples were taken and a 95% confidence interval was computed, 5 of them would be in the interval from 900 to 1100.

Answer by stanbon(75887)

A spinner is divided in half with one half A and the other B. what is the probability that a spinner spun eight times will land on B at least four times?
-------------------
Binomial Problem with n=8 and p= 1/2
---
P(x>=4) = 1-P(0<= x<=3) = 1-binomcdf(8,1/2,3) = 0.6367
=======================
cheers,
Stan H.